OpenStudy (anonymous):

John Gray bought a basic car for $8,250.00, with options that cost $324.00. There's a 6% sales tax in his state and a combined $50.00 license and registration fee. What was John's total cost? A. $9,138.44 B. $10,000.00 C. $14,247.00 D. $9,141.44

OpenStudy (crashonce):

8250+324+50+8574x0.06

OpenStudy (danjs):

you want to total cost of car and options 8250 + 324 = 8574 Plus 6% of that added on, total of 106 percent 1.06*(8574) = 9088.44 plus the 50 dollar fee 9088.44 + 50 = 9138.44 total cost

OpenStudy (anonymous):

\(\rlap{\color{red}{\huge\bigstar}\huge \color{green}{ \text{Welcome to Open Study! }}\color{red}\bigstar}{\; \color{aqua}{\huge\bigstar}\huge \color{aqua}{\text{Welcome to Open Study! }}\color{yellow}\bigstar}\)

OpenStudy (anonymous):

@muneca77